Please help me fill in the blanks.
x^2 + 2/3x + [(2/3)/2]^2 = (x + ____)^2
----
x^2 + (2/3)x + (1/3)^2 = (x+(1/3))^2
=============================================
Cheers,
Stan H.

Answer by stanbon(75887)

Please help me fill in the blanks.
x^2 + 2/3x + [(2/3)/2]^2 = (x + ____)^2
----
x^2 + (2/3)x + (1/3)^2 = (x+(1/3))^2
=============================================
Cheers,
Stan H.